OpenStudy (anonymous):

How do you solve these equations they don’t show me in my book. Once I know how to solve these I can solve the others. (X-81)/(x-3) (X5+y5)/(x+y) (a3+6a2+12a+8)/ (a+2) (5a3+8a2-23a-1)/ (5a2-7a-2)

OpenStudy (anonymous):

first one...(X-81)/(x-3)=0 x=81, 3

OpenStudy (anonymous):

oh there was a divide sign !!!!!!! errrrrrrr

OpenStudy (anonymous):

lets do it again

OpenStudy (anonymous):

(x-9)(x+9)/x-3) (x-3)(x+3)(x+9) /(x-3) (x+3)(x+9)

OpenStudy (anonymous):

ok now third one...

OpenStudy (anonymous):

(a3+6a2+12a+8)/ (a+2) (a+2)(a2+4a+4)/(a+2) (a+2)^2

OpenStudy (anonymous):

now fourth one..

OpenStudy (anonymous):

I think I might have asked the question wrong because and I am confusing you. They want the answer to look something like this (x2 -6x +9)/ (x-3) = (x-3)

OpenStudy (anonymous):

I've done the same too

OpenStudy (anonymous):

not to doubt you but if you doing the same thing give the answer to (x2 +8x-16)/(x+4)

OpenStudy (anonymous):

your question is wrong ..it should have been (x2 +8x+16)/(x+4)= (x+4)

OpenStudy (anonymous):

the answer (x+4) is what I was look for

OpenStudy (anonymous):

thats what I had to solve for

OpenStudy (anonymous):

happy now??

OpenStudy (anonymous):

noo ( again not trying to doubt your abilities) what I was checking was if your answers would look like mine. Now that I see they are what I am asking you to do is tell me the answer to the equations, because I can't see the answers your trying to give.

OpenStudy (anonymous):

Do u want me to do any other ques?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

which one ?

OpenStudy (anonymous):

I'm haveing alot of trouble with these types of questions (X5+y5)/(x+y)

OpenStudy (anonymous):

okay..do u know long division method ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i feel u can get this one done by that..

OpenStudy (anonymous):

I've tried but the answer keeps coming up wrong

OpenStudy (anonymous):

'keeps coming' means u have tried more than once... haven't u ???

OpenStudy (anonymous):

yes, but I get messed up in all the variables

OpenStudy (anonymous):

i m going to do it on paper...I'll write the answer here...lets see what we going to get...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x4-x3y+x2y2-xy3+y3

OpenStudy (anonymous):

is that what u got?

OpenStudy (anonymous):

no, so how did you get there

OpenStudy (radar):

Oldnick, are the X and x the same variable

OpenStudy (anonymous):

ahhh !!! made a lil mistake in the end...ig guess

OpenStudy (anonymous):

no ...i think i was right...@radar whats ur thoughts??

OpenStudy (anonymous):

@radar: yes

OpenStudy (anonymous):

@ankur504: how did you get to your answer?

OpenStudy (anonymous):

long division

OpenStudy (radar):

Is this the problem as written in your text or worksheet?

OpenStudy (anonymous):

@ankur504: yeah I now that give me the proses you used.

OpenStudy (anonymous):

I dont know how to type that method here...sorry

OpenStudy (radar):

\[(x ^{5}+y ^{5})/(x+y)\]

OpenStudy (anonymous):

@ankur504: use the equation tool bar

OpenStudy (anonymous):

wait a min..

OpenStudy (anonymous):

@radar: the first variable looks wrong

OpenStudy (radar):

do you mean it is not x^5???

OpenStudy (radar):

Please restate the problem

OpenStudy (anonymous):

look at this pic..i havent done the full problem but thats enough to give u an idea

OpenStudy (anonymous):

@radar and @oldnick..what u have to say

OpenStudy (anonymous):

@ankur504: I can't really see what your doing

OpenStudy (anonymous):

i m exhausted...sry ...thats all I could do...my bad (

OpenStudy (amistre64):

so whats our equation that we are longjohning?

OpenStudy (radar):

\[(x+y)(x ^{4}-yx ^{3}+y ^{2}x ^{2}-y ^{3}x+y ^{4})\]

OpenStudy (anonymous):

(X5+y5)/(x+y)

OpenStudy (radar):

amistre64 we are trying to do (x^5+y^5/(x+Y)

OpenStudy (anonymous):

@amistre: could you please have a look at the picture which I've uploaded top see if I m correct or not?

OpenStudy (amistre64):

pic looks good, from what I can read :)

OpenStudy (radar):

I think I put the factors up there the (x+y) would cancel and leave the (x^4..............y^4)

OpenStudy (amistre64):

x^4 -x^3 y --------------------- x+y | x^5 + y^5 -x^5 -x^4 y --------------- -x^4 y +y^5 +x^4y +x^3 y^2 ----------------- +x^3 y^2 +y^5 and it keeps going till you get no x values to plug into it

OpenStudy (anonymous):

@amistre64: I can't understand what you did there.

OpenStudy (amistre64):

its just long division.... the same way you do numbers....but with xs and ys

OpenStudy (amistre64):

les try 27/5 we want a number that is less than but close to 5*x = 27 lets use 5 5. <- when we multiply this to the 5 we "subtract" it from the 27 ------ 5 | 27.00000 - 25 <- 5*5 = 25....but!! 27-25 is what we want for an answer ----- 20 <- this is whats left over and drop down the next "number"

OpenStudy (radar):

oldnick, i suggest you factor the (x^5+y^5) first using the rule \[x ^{n}+y ^{n}\] when n is odd. n is 5 so it is odd.

OpenStudy (anonymous):

thats the problem is that the book gives an example on how to do the equation with (x3 +1)/(x+1). but not with only variables, so I don't now what to do with variables and large numbers.

OpenStudy (amistre64):

5.4 ------ 5 | 27.00000 - 25 ----- 20 <- now we want a number that is 5*x = 20, use 4 :) -20 <- 5*4 = 20...but!! we want 20 - 20 for an answer.. ------- 0

OpenStudy (anonymous):

@radar: whats the xn+yn rule

OpenStudy (amistre64):

the variables that cancel go away...and you add the ones that are alike and bring down the ones that arent....

OpenStudy (radar):

\[(x+y)(x ^{n-1)}-yx ^{n-2}+a ^{2}x ^{n-3}-...+y ^{n-1})\] you can use this then as you see the x+y divides into the x=y giving 1 and you have the long one left which is the answer4

OpenStudy (amistre64):

(x) * ? is close to x^5. lets use x^4 (x+y)x^4 = x^5 + x^4 y so we subtract that to get a new round -(x^5 +x^4 y) = -x^5 -x^4 y x^4 --------------------- x+y | x^5 + y^5 -x^5 -x^4 y <- we subtract it --------------- 0 -x^4 y +y^5 <- the x^5 goes to zero and the only things left that didnt subtract to zero is brought down and worked again

OpenStudy (radar):

Corrections; change a to a y. remember in this case n=5

OpenStudy (amistre64):

(x) * ? is close to -x^4 y. lets use -x^3 y (x+y)-x^3 y = -x^4 y -x^3 y^2 so we subtract that to get a new round -(-x^4 y -x^3 y^2) = x^4 y +x^3 y^2 x^4 -x^3 y --------------------- x+y | x^5 + y^5 -x^5 -x^4 y --------------- -x^4 y +y^5 x^4 y +x^3 y^2 <- subtract it ----------------- 0 + x^3 y^2 +y^5 <-whats left that didnt subtract to zero is brought down and worked again

OpenStudy (amistre64):

thats the easiest way I know to explain it...

OpenStudy (amistre64):

on a keyboard that is :)

OpenStudy (radar):

The keyboard does provide limitation

OpenStudy (anonymous):

I'm not sure if this will help but this is what they want me to do.

OpenStudy (amistre64):

and that is exactly what I showed you ;)

OpenStudy (radar):

Then the amistre64 method (long division) is the way to go

OpenStudy (anonymous):

heres the problem what you showed me I can't understand

OpenStudy (amistre64):

x^5 + y^5 is the same as: x^5 +0x^4y +0x^3 y^2+0x^2 y^3+0x y^4+y^5 if you expand it out...

OpenStudy (anonymous):

right I understand that

OpenStudy (amistre64):

that is just impractical on the keyboard to try to use...

OpenStudy (amistre64):

so instead of using all those 0 place holders... ignore them.

OpenStudy (amistre64):

add together your like terms...and drop the rest down

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

but there are no like terms

OpenStudy (amistre64):

x^5 +0x^4 y + ............+ y^5 -x^5 -x^4 y --------------- -x^4 y + ............. +y^5 right?

OpenStudy (anonymous):

thats where you loss me

OpenStudy (amistre64):

I just included the "like term" for you to see that its just a place holder

OpenStudy (radar):

The sum of equal powers can be factored. different rules for even vs odd powers. Not really different rules, just different patterns.

OpenStudy (amistre64):

y^5 -x^4 y ------- -x^4 y + y^5 thats all there is to it....

OpenStudy (amistre64):

whatever doesn subtract from each other...carries on down ....... do you see that?

OpenStudy (amistre64):

like terms dont combine...so both terms drop down as they are..

OpenStudy (anonymous):

what do the ------ stand for?

OpenStudy (amistre64):

unlike terms dont combine is what it should read... my fingers got lazy on me :)

OpenStudy (amistre64):

thats an equals bar underneath the terms to be added. 4 +5 ---- 9

OpenStudy (amistre64):

we can add across like this: 4+5 = 9 or we can add up and down like this: 4 +5 ---- 9

OpenStudy (amistre64):

when we do long division; we add/subtract in the vertical way.... keeps everything tidy

OpenStudy (amistre64):

yes?no? makes sense? doesnt? i need to hear from you...